The information below pertains to Ivanhoe Company for 2021.

Net income for the year
$1,170,000
8% convertible bonds issued at par ($1,000 per bond); each bond is convertible into
    30 shares of common stock

2,060,000
6% convertible, cumulative preferred stock, $100 par value; each share is convertible
     into 3 shares of common stock

3,990,000
Common stock, $10 par value
6,290,000
Tax rate for 2021
20%
Average market price of common stock
$25 per share

There were no changes during 2021 in the number of common shares, preferred shares, or convertible bonds outstanding. There is no treasury stock. The company also has common stock options (granted in a prior year) to purchase 82,400 shares of common stock at $20 per share.

(a) Compute basic earnings per share for 2021. (Round answer to 2 decimal places, e.g. $2.55.)

(b) Compute diluted earnings per share for 2021. (Round answer to 2 decimal places, e.g. $2.55.)
Answer #1

(a) Basis EPS=(Net income-Preferred dividend)/Weight average equity shares outstanding

Preferred dividend=(3,990,000*6%)=$239,400)

Weighted average number of equity shares=6,290,000/10=629,000 shares

Basic EPS=($1,170,000-$239,400)/629,000

Basic EPS=$930,600/62,900

Basic EPS=$1.48 per share


(b) Diluted EPS=(Net income-Preferred dividend)+Interest savings(net of tax)/(Weighted average number of equity shares+Potentially dilutive common shares)

Diluted EPS=(($1,170,000-$239,400)+$131,840)/ (629,000+16,480+61,800)

Diluted EPS=$1,062,440/707,280

Diluted EPS=$1.50 per share

Working notes:

1.Interest savings(Net of tax)=$2,060,000*8%*(1-20%)

= $131,840


2.(a)Potentially dilutive common shares
Incremental shares=(Market price-Option price/Market price )*Number of options

Incremental shares=(($25-$20)/$25)*82,400 shares

Incremental shares=16,480 shares


2.(b).($2,060,000/1,000)*30 shares per bond
=61,800 shares


answered by: Allen